OpenStudy (anonymous):

Simplify. (2a – 5)2 A. 42a – 7 + 16 B. 4a2 – 25 C. 4a2– 20a + 10 D. 4a2 – 20a + 25

OpenStudy (anonymous):

foil?

OpenStudy (anonymous):

(2a−5)(2) =(2a+(−5))(2) =(2a)(2)+(−5)(2) =4a−10

OpenStudy (anonymous):

(2a – 5)^2 (2x-5)(2x-5) 4x^-10x-10x+25 4x^2-20x+25

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

So I think its C or D

OpenStudy (anonymous):

Thank you guys!

OpenStudy (anonymous):

Yeah, I think D